How to calculate the energy spent on heating a metallic and plastic glucose barrel up to 50 ° C
1
Expert's answer
2020-07-29T09:43:11-0400

Let's assume, the glucose barrel has the mass "m = 300 kg" and was kept at temperature "T_0 = 25\degree C = 298 K". The specific heat capacity of glucose is "с = 1255 \dfrac{J}{kg\cdot K}".

Then the energy spent on heating up to "T = 50\degree C = 323 K" will be:


"Q = mc(T - T_0) = 300\cdot 1255\cdot (323 - 298) \approx 9.4MJ"

Answer. 9.4 MJ.